A Complete Bibliography of Internet FYI Documents

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

19 July 2005 Version 1.13

Title word cross-reference

$$$$$ [38].

1 [1]. 10 [10]. 11 [11]. 12 [12, 26]. 13 [13]. 14 [14]. 15 [15]. 16 [16]. 17 [17]. 18 [18]. 19 [19].

2 [2]. 20 [20]. 21 [21]. 22 [22]. 23 [23]. 24 [24]. 25 [25]. 26 [26]. 27 [27]. 28 [28]. 29 [29].

3 [3]. 30 [30]. 31 [31]. 32 [32]. 33 [33]. 34 [34]. 35 [35]. 36 [36]. 37 [37]. 38 [38].

4 [4].

5 [5].

6 [6]. 7 [7]. 8 [8]. 9 [9]. Accuracy [15]. Administrative [37]. Advanced [21]. Advertise [38]. Anonymous [24]. Answers [4, 7]. Anticipate [16]. Arts [31]. Asked [4, 22, 7]. Bibliography [19, 3]. Biographies [9]. Building [12]. Catalog [11, 2]. Catalog-96 [11]. Catalogue [29]. Center [15, 31]. Choosing [5]. Commonly [4, 7]. computer [5]. Connecting [16]. Databases [15]. Debugging [2, 27]. Devices [2]. Directory [13, 14]. DNS [27]. DON’T [35]. E-Mail [38]. ENEMIES [38]. Engineering [17]. Executive [13]. experienced [7]. FAST [38]. Force [17]. Frequently [22]. FTP [24]. FYI [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. Glossary [18, 36]. Gold [10]. Groups [25]. Guide [17, 23, 37]. Guidelines [26, 28, 35]. Handbook [8, 34]. Hobbes [32]. Humanities [31]. IAB [9]. IESG [9]. IETF [17]. Implementations [11]. Implementors [33]. Information [12, 15, 25, 3]. Infrastructure [12, 37]. Institutions [16]. Interconnected [2]. Internet [4, 7, 20, 9, 16, 17, 18, 19, 30, 31, 32, 36, 37]. Internets [2]. Internetworking [19, 26, 3]. Introducing [19]. Introduction [1, 13]. Introductory [19]. IRSG [9]. Issues [15]. K-12 [26]. Mail [38]. Mailings [35]. MAKE [38]. management [2]. Mass [35]. Materials [29]. Members [9]. Monitoring [2]. name [5]. Netiquette [28]. Network [12, 15, 23, 29, 2]. Networked [25].

Networks [10]. Newsgroups [38]. Notes [1]. Novice [17].

Overview [14].

IP [2, 30]. Places [10]. Postings [35]. Primer [30]. Privacy [15]. Procedures [37]. Protocol [13, 14].

Questions [4, 7, 22].

Readings [19]. Report [25]. Resource [23]. Responsibly [38]. Retrieval [25].

Schools [22]. Searching [10]. Security [8, 34, 36]. Services [12, 13, 14]. Set [35]. Sharing [31]. Short [19]. Should [16]. Site [8]. Some [33]. spam* [35]. SPEW [35]. Stage [31]. start [3]. Status [25]. Survey [21]. System [6].

Tao [17]. Task [17]. TCP [2, 30, 33]. TCP/IP [2, 30]. Technical [14]. Testing [33]. thar [10]. them [10]. There [10]. Timeline [32]. tool [2]. Tools [2, 23, 25, 27, 30, 33]. Training [29]. Treasure [10].

Unsolicited [35]. Usages [21]. Use [24]. User [4, 18, 34, 7]. Using [13, 14, 38]. Utilities [30].

where [3]. Who [9]. Window [6]. Wrong [10].

X [6]. X.500 [11, 13, 14, 21].

References

FYI0001

[1] G. S. Malkin and J. K. Reynolds. FYI 1: FYI on FYI: Introduction to the FYI notes, March 1, 1990. URL ftp://ftp.isi.edu/in-

    notes/fyi/fyi1.txt,   ftp://ftp.isi.edu/in-notes/fyi/rfc1150.
    txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi1.txt,   ftp://ftp.
    math.utah.edu/pub/rfc/rfc1150.txt. See also RFC1150 [?].

FYI0002

[2] R. Enger and J. Reynolds. FYI 2: FYI on a network management tool catalog: Tools for monitoring and debugging TCP/IP Internets and interconnected devices, June 1993. URL ftp://ftp.isi.edu/in-notes/

    fyi/fyi2.txt,   ftp://ftp.isi.edu/in-notes/fyi/rfc1147.txt,

ftp://ftp.isi.edu/in-notes/fyi/rfc1470.txt,   ftp://ftp.math.
utah.edu/pub/rfc/fyi2.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1147.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1470.txt. See

also RFC1470 [?]. Obsoletes RFC1147 [?].

FYI0003

[3] K. L. Bowers, T. L. LaQuey, J. K. Reynolds, K. Roubicek, M. K. Stahl, and A. Yuan. FYI 3: FYI on where to start: A bibliography of internetworking information, August 1, 1990. URL ftp://ftp.isi.edu/in-

notes/fyi/fyi3.txt,   ftp://ftp.isi.edu/in-notes/fyi/rfc1175.
txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi3.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1175.txt. See also RFC1175 [?].

FYI0004

[4] R. Plzak, A. Wells, and E. Krol. FYI 4: FYI on questions and answers — answers to commonly asked “New Internet User” questions, August

1999. URL ftp://ftp.isi.edu/in-notes/fyi/fyi4.txt,   ftp://
ftp.isi.edu/in-notes/fyi/rfc1594.txt,   ftp://ftp.isi.edu/in-
notes/fyi/rfc2664.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi4.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1594.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc2664.txt. See also RFC2664 [?]. Obsoletes

RFC1594 [?].

FYI0005

[5] D. Libes. FYI 5: Choosing a name for your computer, Au-

gust 1, 1990. URL ftp://ftp.isi.edu/in-notes/fyi/fyi5.txt,
ftp://ftp.isi.edu/in-notes/fyi/rfc1178.txt,   ftp://ftp.math.
utah.edu/pub/rfc/fyi5.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1178.txt. See also RFC1178 [?].

FYI0006

[6] R. W. Scheifler. FYI 6: FYI on the X Window System, Jan-

uary 1, 1991. URL ftp://ftp.isi.edu/in-notes/fyi/fyi6.txt,
ftp://ftp.isi.edu/in-notes/fyi/rfc1198.txt,   ftp://ftp.math.
utah.edu/pub/rfc/fyi6.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1198.txt. See also RFC1198 [?].

FYI0007

[7] G. S. Malkin, A. N. Marine, and J. K. Reynolds. FYI 7: FYI on questions and answers: Answers to commonly asked “experienced Internet user” questions, February 1, 1991. URL ftp://ftp.isi.edu/in-

notes/fyi/fyi7.txt,   ftp://ftp.isi.edu/in-notes/fyi/rfc1207.

txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi7.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1207.txt. See also RFC1207 [?].

FYI0008

[8] B. Fraser. FYI 8: Site security handbook, September 1997. URL

ftp://ftp.isi.edu/in-notes/fyi/fyi8.txt,   ftp://ftp.isi.edu/
in-notes/fyi/rfc1244.txt,   ftp://ftp.isi.edu/in-notes/fyi/rfc2196.
txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi8.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1244.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2196.txt. See also RFC2196 [?]. Obsoletes RFC1244 [?].

FYI0009

[9] G. Malkin. FYI 9: Who’s who in the Internet: Biographies of IAB, IESG

and IRSG members, May 1992. URL ftp://ftp.isi.edu/in-notes/
fyi/fyi9.txt,   ftp://ftp.isi.edu/in-notes/fyi/rfc1251.txt,
ftp://ftp.isi.edu/in-notes/fyi/rfc1336.txt,   ftp://ftp.math.
utah.edu/pub/rfc/fyi9.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1251.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1336.txt. See

also RFC1336 [?]. Obsoletes RFC1251 [?].

FYI0010

[10] J. Martin. FYI 10: There’s gold in them thar networks! or searching for treasure in all the wrong places, January 1993. URL ftp://ftp.isi.

edu/in-notes/fyi/fyi10.txt,   ftp://ftp.isi.edu/in-notes/fyi/
rfc1290.txt,   ftp://ftp.isi.edu/in-notes/fyi/rfc1402.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/fyi10.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1290.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1402.

txt. See also RFC1402 [?]. Obsoletes RFC1290 [?].

FYI0011

[11] C. Apple and K. Rossen. FYI 11: X.500 implementations catalog-96, April

1997. URL ftp://ftp.isi.edu/in-notes/fyi/fyi11.txt,   ftp://
ftp.isi.edu/in-notes/fyi/rfc1632.txt,   ftp://ftp.isi.edu/in-
notes/fyi/rfc2116.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi11.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1632.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc2116.txt. See also RFC2116 [?]. Obsoletes

RFC1632 [?].

FYI0012

[12] D. Sitzler, P. Smith, and A. Marine. FYI 12: Building a network information services infrastructure, February 1992. URL ftp://ftp.

isi.edu/in-notes/fyi/fyi12.txt,   ftp://ftp.isi.edu/in-notes/
fyi/rfc1302.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi12.txt,

 ftp://ftp.math.utah.edu/pub/rfc/rfc1302.txt. See also RFC1302

[?].

FYI0013

[13] C. Weider and J. Reynolds. FYI 13: Executive introduction to directory services using the X.500 protocol, March 1992. URL ftp://ftp.

isi.edu/in-notes/fyi/fyi13.txt,   ftp://ftp.isi.edu/in-notes/
fyi/rfc1308.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi13.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1308.txt. See also RFC1308

[?].

FYI0014

[14] C. Weider, J. Reynolds, and S. Heker. FYI 14: Technical overview of directory services using the X.500 protocol, March 1992. URL ftp://ftp.

isi.edu/in-notes/fyi/fyi14.txt,   ftp://ftp.isi.edu/in-notes/
fyi/rfc1309.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi14.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1309.txt. See also RFC1309

[?].

FYI0015

[15] J. Curran and A. Marine. FYI 15: Privacy and accuracy issues in network information center databases, August 1992. URL ftp://ftp.

isi.edu/in-notes/fyi/fyi15.txt,   ftp://ftp.isi.edu/in-notes/
fyi/rfc1355.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi15.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1355.txt. See also RFC1355

[?].

FYI0016

[16] ACM SIGUCCS. FYI 16: Connecting to the Internet — what connecting institutions should anticipate, August 1992. URL ftp://ftp.

isi.edu/in-notes/fyi/fyi16.txt,   ftp://ftp.isi.edu/in-notes/
fyi/rfc1359.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi16.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1359.txt. See also RFC1359

[?].

FYI0017

[17] S. Harris. FYI 17: The tao of IETF — A novice’s guide to the Internet Engineering Task Force, August 2001. URL ftp://ftp.isi.

edu/in-notes/fyi/fyi17.txt,   ftp://ftp.isi.edu/in-notes/fyi/
rfc1718.txt,   ftp://ftp.isi.edu/in-notes/fyi/rfc3160.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/fyi17.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1718.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc3160.

txt. See also RFC3160 [?]. Obsoletes RFC1718 [?].

FYI0018

[18] G. Malkin. FYI 18: Internet users’ glossary, August 1996. URL ftp://ftp.

isi.edu/in-notes/fyi/fyi18.txt,   ftp://ftp.isi.edu/in-notes/
fyi/rfc1392.txt,   ftp://ftp.isi.edu/in-notes/fyi/rfc1983.txt,
 ftp://ftp.math.utah.edu/pub/rfc/fyi18.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1392.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1983.txt. See also RFC1983 [?]. Obsoletes RFC1392 [?].

FYI0019

[19] E. Hoffman and L. Jackson. FYI 19: FYI on introducing the Internet — A short bibliography of introductory internetworking readings, May

1993. URL ftp://ftp.isi.edu/in-notes/fyi/fyi19.txt,   ftp:/
/ftp.isi.edu/in-notes/fyi/rfc1463.txt,   ftp://ftp.math.utah.
edu/pub/rfc/fyi19.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1463.

txt. See also RFC1463 [?].

FYI0020

[20] E. Krol and E. Hoffman. FYI 20: FYI on “What is the Internet?”, May

1993. URL ftp://ftp.isi.edu/in-notes/fyi/fyi20.txt,   ftp:/
/ftp.isi.edu/in-notes/fyi/rfc1462.txt,   ftp://ftp.math.utah.
edu/pub/rfc/fyi20.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1462.

txt. See also RFC1462 [?].

FYI0021

[21] C. Weider and R. Wright. FYI 21: A survey of advanced usages of X.500,

July 1993. URL ftp://ftp.isi.edu/in-notes/fyi/fyi21.txt,
ftp://ftp.isi.edu/in-notes/fyi/rfc1491.txt,   ftp://ftp.math.
utah.edu/pub/rfc/fyi21.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1491.txt. See also RFC1491 [?].

FYI0022

[22] J. Sellers and J. Robichaux. FYI 22: Frequently asked questions for

schools, May 1996. URL ftp://ftp.isi.edu/in-notes/fyi/fyi22.txt,
 ftp://ftp.isi.edu/in-notes/fyi/rfc1578.txt,   ftp://ftp.isi.
edu/in-notes/fyi/rfc1941.txt,   ftp://ftp.math.utah.edu/pub/rfc/
fyi22.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1578.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1941.txt. See also RFC1941 [?]. Ob-

soletes RFC1578 [?].

FYI0023

[23] EARN Staff. FYI 23: Guide to network resource tools, March

1994. URL ftp://ftp.isi.edu/in-notes/fyi/fyi23.txt,   ftp:/

/ftp.isi.edu/in-notes/fyi/rfc1580.txt,   ftp://ftp.math.utah.
edu/pub/rfc/fyi23.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1580.

txt. See also RFC1580 [?].

FYI0024

[24] P. Deutsch, A. Emtage, and A. Marine. FYI 24: How to use anonymous

FTP, May 1994. URL ftp://ftp.isi.edu/in-notes/fyi/fyi24.txt,
ftp://ftp.isi.edu/in-notes/fyi/rfc1635.txt,   ftp://ftp.math.
utah.edu/pub/rfc/fyi24.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1635.txt. See also RFC1635 [?].

FYI0025

[25] J. Foster. FYI 25: A status report on networked information retrieval: Tools and groups, August 1994. URL ftp://ftp.isi.

edu/in-notes/fyi/fyi25.txt,   ftp://ftp.isi.edu/in-notes/fyi/
rfc1689.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi25.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1689.txt. See also RFC1689 [?].

FYI0026

[26] J. Gargano and D. Wasley. FYI 26: K-12 internetworking guidelines,

November 1994. URL ftp://ftp.isi.edu/in-notes/fyi/fyi26.txt,
ftp://ftp.isi.edu/in-notes/fyi/rfc1709.txt,   ftp://ftp.math.
utah.edu/pub/rfc/fyi26.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1709.txt. See also RFC1709 [?].

FYI0027

[27] A. Romao. FYI 27: Tools for DNS debugging, November 1994. URL ftp:

//ftp.isi.edu/in-notes/fyi/fyi27.txt,   ftp://ftp.isi.edu/in-
notes/fyi/rfc1713.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi27.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1713.txt. See also

RFC1713 [?].

FYI0028

[28] S. Hambridge. FYI 28: Netiquette guidelines, October 1995. URL ftp:

//ftp.isi.edu/in-notes/fyi/fyi28.txt,   ftp://ftp.isi.edu/in-
notes/fyi/rfc1855.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi28.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1855.txt. See also

RFC1855 [?].

FYI0029

[29] J. Foster, M. Isaacs, and M. Prior. FYI 29: Catalogue of network train-

ing materials, October 1996. URL ftp://ftp.isi.edu/in-notes/fyi/

fyi29.txt,   ftp://ftp.isi.edu/in-notes/fyi/rfc2007.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/fyi29.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc2007.txt. See also RFC2007 [?].

FYI0030

[30] G. Kessler and S. Shepard. FYI 30: A primer on Internet and TCP/IP tools

and utilities, June 1997. URL ftp://ftp.isi.edu/in-notes/fyi/fyi30.
txt,   ftp://ftp.isi.edu/in-notes/fyi/rfc1739.txt,   ftp://ftp.
isi.edu/in-notes/fyi/rfc2151.txt,   ftp://ftp.math.utah.edu/pub/
rfc/fyi30.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1739.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2151.txt. See also RFC2151

[?]. Obsoletes RFC1739 [?].

FYI0031

[31] J. Max and W. Stickle. FYI 31: Humanities and arts: Sharing center stage on the Internet, October 1997. URL ftp://ftp.

isi.edu/in-notes/fyi/fyi31.txt,   ftp://ftp.isi.edu/in-notes/
fyi/rfc2150.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi31.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2150.txt. See also RFC2150

[?].

FYI0032

[32] R. Zakon. FYI 32: Hobbes’ Internet timeline, November 1997. URL ftp:

//ftp.isi.edu/in-notes/fyi/fyi32.txt,   ftp://ftp.isi.edu/in-
notes/fyi/rfc2235.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi32.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2235.txt. See also

RFC2235 [?].

FYI0033

[33] S. Parker and C. Schmechel. FYI 33: Some testing tools for TCP implemen-

tors, August 1998. URL ftp://ftp.isi.edu/in-notes/fyi/fyi33.txt,
 ftp://ftp.isi.edu/in-notes/fyi/rfc2398.txt,   ftp://ftp.math.
utah.edu/pub/rfc/fyi33.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc2398.txt. See also RFC2398 [?].

FYI0034

[34] E. Guttman, L. Leong, and G. Malkin. FYI 34: Users’ security handbook,

February 1999. URL ftp://ftp.isi.edu/in-notes/fyi/fyi34.txt,
ftp://ftp.isi.edu/in-notes/fyi/rfc2504.txt,   ftp://ftp.math.
utah.edu/pub/rfc/fyi34.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc2504.txt. See also RFC2504 [?].

FYI0035

[35] S. Hambridge and A. Lunde. FYI 35: DON’T SPEW: A set of guidelines for mass unsolicited mailings and postings (spam*), June

1999. URL ftp://ftp.isi.edu/in-notes/fyi/fyi35.txt,   ftp:/
/ftp.isi.edu/in-notes/fyi/rfc2635.txt,   ftp://ftp.math.utah.
edu/pub/rfc/fyi35.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2635.

txt. See also RFC2635 [?].

FYI0036

[36] R. Shirey. FYI 36: Internet security glossary, May 2000. URL ftp://ftp.

isi.edu/in-notes/fyi/fyi36.txt,   ftp://ftp.isi.edu/in-notes/
fyi/rfc2828.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi36.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2828.txt. See also RFC2828

[?].

FYI0037

[37] Z. Wenzel, J. Klensin, R. Bush, and S. Huter. FYI 37: Guide to administrative procedures of the Internet infrastructure, August

2000. URL ftp://ftp.isi.edu/in-notes/fyi/fyi37.txt,   ftp:/
/ftp.isi.edu/in-notes/fyi/rfc2901.txt,   ftp://ftp.math.utah.
edu/pub/rfc/fyi37.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2901.

txt. See also RFC2901 [?].

FYI0038

[38] E. Gavin, D. Eastlake 3rd, and S. Hambridge. FYI 38: How to advertise responsibly using E-mail and newsgroups or — how NOT to $$$$$ MAKE ENEMIES FAST! $$$$$, April 2001. URL ftp://ftp.

isi.edu/in-notes/fyi/fyi38.txt,   ftp://ftp.isi.edu/in-notes/
fyi/rfc3098.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi38.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc3098.txt. See also RFC3098

[?].